A farmer has 800 ft of fencing material to enclose a rectangular field. The
width of the field is 175 ft. How do I find the length and area of the field?

Answer by checkley77(12844)

2w+2l=800
2*175+2l=800
350+2l=800
2l=800-350
2l=450
l=450/2
l=225 ft. for the length.
Proof:
2*175+2*225=800
350+450=800
800=800
area=175*225=39,375 ft^2.